# Orbifold modifications of complex analytic varieties

### v4

@@ Abstract.
We prove that if $X$ is a compact complex analytic variety, which has quotient singularities in codimension 2, then there is a projective bimeromorphic morphism $f\colon Y\to X$, such that $Y$ has quotient singularities, and that the indeterminacy locus of $f^{-1}$ has codimension at least 3 in $X$. As an application, we deduce the Bogomolov-Gieseker inequality on orbifold Chern classes for stable reflexive coherent sheaves on compact Kähler varieties which have quotient singularities in codimension 2.
@@ 1. Introduction
Let $X$ be a compact complex analytic variety. Assume that $X$ has quotient singularities in codimension 2. In other words, there is a closed analytic subset $V$ of codimension at least 3, such that $X\setminus V$ has quotient singularities. Then there is a projective bimeromorphic morphism $f\colon Y\to X$ such that $Y$ has quotient singularities, and that the indeterminacy locus of $f^{-1}$ has codimension at least 3 in $X$.
By combining the theorem with the existence of orbifold Hermitian-Einstein metrics in [Fau22], and by using a continuity argument, we deduce the following Bogomolov-Gieseker inequality.
Theorem 1.3.
Let $(X,\omega)$ be a compact Kähler variety of dimension $n\geq 2$, which has quotient singularities in codimension 2. Let $\mathcal{F}$ be a $\omega$-stable reflexive coherent sheaf of rank $r\geq 2$ on $X$. Then we have the following inequality on orbifold Chern classes of $\mathcal{F}$,
$\Big{(}2r\hat{c}_{2}(\mathcal{F})-(r-1)\hat{c}_{1}(\mathcal{F})^{2}\Big{)}\cdot[\omega]^{n-2}\geq 0.$
For the precise definition of the Chern classes $\hat{c}$, we refer to Section 9. We refer to [GP24] for a pure analytic approach to Bogomolov-Gieseker inequalities when $X$ is a klt threefold. The inequality of Theorem 1.3 plays an important role in the abundance theorem for compact Kähler threefolds (see [CHP23], [DO23] and [GP24]). We also remark that it is stronger than the one obtained by taking a desingularization of $X$. For example, when $X$ is a projective surface with quotient singularities, such a comparison was shown in [Lan00, Proposition 2.4.(6) and Theorem 5.1].
We note that when $X$ is projective, a stronger version of Theorem 1.2 was proved by Xu in [LT19]. The proof approximately proceeds as follows. We let $X^{\circ}$ be the largest open subset on which $X$ has quotient singularities. By considering the frame bundle on the local orbifold charts, we can obtain a smooth quasi-projective variety $Y^{\circ}$, such that $X^{\circ}=Y^{\circ}/G$, where $G=\mathrm{GL}_{n}(\mathbb{C})$ and $n$ is the dimension of $X$ (see [Kre09]). The action of $G$ on $Y^{\circ}$ only has finite stabilizers. There exists a $G$-equivariant projective compactification ${Y}$ of $Y^{\circ}$, such that $G$ acts linearly on ${Y}$. Then up to applying Kirwan’s partial resolution in [Kir85], one may assume that the strictly semi-stable locus in ${Y}$ is empty. Then the GIT quotient ${Y}\sslash G$ is a projective compactification of $X^{\circ}$ which has quotient singularities only. Finally, by using equivariant resolution of singularities to the birational map ${Y}\sslash G\dashrightarrow X$, we can obtain a projective compactification $Z$ of $X^{\circ}$ such that $Z$ has quotient singularities and the natural birational map $Z\dashrightarrow X$ is a morphism. The main difficulty to adapt this method in the setting of general complex analytic varieties is that, the theory of GIT relies essentially on the existence of an ample line bundle, which is not the case if $X$ is not projective.
@@ 2. Preliminaries
We fix some notation and prove some elementary results in this section. Throughout this paper, the symbol $\mathbb{D}$ stands for a disc contained in $\mathbb{C}$ centered at 00. We will write $\mathbb{D}_{r}$ if we need to stress the radius $r$ of $\mathbb{D}$. We denote the origin of $\mathbb{C}^{n}$ by $\mathbf{0}_{n}$. By a polydisc of dimension $n$, we refer to $\mathbb{D}^{n}$, which is always assumed centered at the origin of $\mathbb{C}^{n}$.
@@ 2.A. Complex analytic varieties and their singularities
A complex analytic variety $X$ is a reduced and irreducible complex analytic space. We will denote by $X_{\operatorname{\textsubscript{\rm sm}}}$ its smooth locus and by $X_{\operatorname{\textsubscript{\rm sing}}}$ its singular locus. A smooth complex analytic variety is also called a complex manifold.
@@
We may assume hat $U=\mathbb{D}^{n-2}\times V$. Then the subset $S\subseteq U$ is identified with $\mathbb{D}^{n-2}\times\{o\}$. Let $\mu\colon\widetilde{V}\to V$ be the minimal resolution, and let $\rho\colon\widetilde{U}\to U$ be the desingularization with $\widetilde{U}=\mathbb{D}^{n-2}\times\widetilde{V}$. Let $\widetilde{T}=\rho^{-1}(T)$, and let $\tilde{x}\in\widetilde{T}$ be a point lying over $x$. With the product structure $\widetilde{U}=\mathbb{D}^{n-2}\times\widetilde{V}$, we can write $\tilde{x}=(\mathbf{0}_{n-2},\tilde{v})$, where $\tilde{v}\in\widetilde{V}$ is contained in the exceptional divisor of $\mu$. Let $\sigma\colon S\to\widetilde{U}$ be a section, which is the composition of the following sequence,
$S=\mathbb{D}^{n-2}\times\{o\}\to\mathbb{D}^{n-2}\times\{\tilde{v}\}\hookrightarrow\widetilde{U}.$
Since $T$ meets $S$ transversally at $x$, we deduce that, at $\tilde{x}$, the Zariski tangent space of $\widetilde{T}$ and the Zariski tangent space of $\sigma(S)$, both viewed as the subspace of the Zariski tangent space of $\widetilde{U}$, intersect transversally. Since $\operatorname{dim}\,\widetilde{T}+\operatorname{dim}\,\sigma(S)=\operatorname{dim}\,\widetilde{U}$, and since $\widetilde{U}$ is smooth, we deduce that $\widetilde{T}$ is smooth at $\tilde{x}$. Therefore, we obtain that $\widetilde{T}$ is smooth. Furthermore, by the adjunction formula, we see that $\widetilde{T}\to T$ is the minimal resolution. This implies that $(x\in T)$ is isomorphic to $(o\in V)$ as surface singularities. ∎
@@
where $p,q$ are holomorphic functions whose terms have degrees at least 2 and 3 respectively. Replacing $z$ by $z-p(y,z)$, we can assume that
@@
Let $f\colon Y\to X$ be a proper bimeromorphic morphism between complex analytic spaces. Since we do not assume that $Y$ is normal, we say that there is a $f$-ample and $f$-exceptional divisor $-H$, if $H$ is a closed analytic subspace of $Y$ contained in the $f$-exceptional locus, such that the ideal sheaf $\mathcal{I}$ of $H$ is a $f$-ample invertible sheaf (see [Nak87, Definition 1.1]). In the proof of the main theorem, we apply several local constructions around a component of the singular locus. The following lemma enables us to extend local modifications to the whole variety.
@@
Let $X$ be a complex analytic space and let $U\subseteq X$ be an open subset. Assume that there is a proper bimeromorpihc morphism $p\colon U^{\prime}\to U$ such that the indeterminacy locus $Z$ of $p^{-1}$ is compact. Then $p$ extends to some proper bimeromorpihc morphism $f\colon X^{\prime}\to X$, which is an isomorphism over $X\setminus Z$. Assume further that there is some $p$-exceptional divisor in $U^{\prime}$ which is $p$-ample over $U$, then $f$ is projective over $X$.
@@
Let $W=X\setminus Z$. Since $p$ is an isomorphism over $U\cap W$, we deduce that $U^{\prime}$ and $W$ glue together along $U\cap W$ to a complex analytic space $X^{\prime}$, and that $p$ extends to a proper bimeromorphic morphism $f\colon X^{\prime}\to X$. If there is a $p$-exceptional and $p$-ample divisor $-H$, then $H$ is compact since $Z$ is compact. Thus $H$ is naturally a closed analytic subspace of $X^{\prime}$. We note that its ideal sheaf in $X$ is also invertible and $f$-ample over $X$. Hence $f\colon X^{\prime}\to X$ is projective. ∎
@@ 2.B. Finite morphisms between complex analytic varieties
In this subsection, we summarize some results concerning finite morphisms between complex analytic varieties. We recall from [GR84, page 47] that a morphism $f\colon Y\to X$ is called finite if it is closed and if every fiber is a finite set. In this case, the critical locus of $f$ is the smallest closed subset $Z\subseteq X$, such that $f$ is étale over $X\setminus Z$, that is, $f$ is locally biholomorphic on $Y\setminus f^{-1}(Z)$. By the divisorial critical locus of $f$, we refer to the codimension 1 part of $Z$. The morphism $f$ is said to be quasi-étale if its critical locus has codimension at least $2$. By abuse of notation, we say that $f$ is Galois of Galois group $G$ if its restriction over $X\setminus Z$ is Galois, and if there is an action of $G$ on $Y$, extending the Galois action of $G$ on $f^{-1}(X\setminus Z)$, such that the quotient $Y/G$ is canonically isomorphic to $X$. We refer to [Car57, Théorème 4] for the quotient of a complex analytic space by a finite group. The following theorem is due to Grauert-Remmert.
@@
Let $X$ be a complex analytic variety, and let $X^{\circ}\subseteq X$ be a dense Zariski open subset. Assume that $X^{\circ}$ is normal and we have a finite étale morphism $p\colon Y^{\circ}\to X^{\circ}$. Then $p$ extends to a finite morphism $p\colon Y\to X$ with $Y$ normal, which is unique up to isomorphism.
@@
Let $r\colon X^{\prime}\to X$ be the normalization. By assumption, $r$ is an isomorphism over $X^{\circ}$. Hence up to replacing $X$ by $X^{\prime}$, we may assume that $X$ is normal. In this case, the theorem is proved in [GR71, Théorème XII.5.4]. ∎
@@
[…]
@@
With the notation above, let $x\in X$ be a point and let $s^{\prime}_{1},...,s^{\prime}_{m}$ be holomorphic functions on $X$. Assume that there are unit functions $\mu_{1},...,\mu_{m}$ such that $s^{\prime}_{i}=\mu_{i}\cdot s_{i}$ for all $i=1,...,m$. Then, up to shrinking $X$ around $x$, there is an isomorphism over $X$
[…]
@@
[…]
@@
[…]
@@
[…]
@@
For the item (1), we first prove the statement on the ramification indices of $p$. Since $X$ is normal and $D_{i}$ is reduced, we deduce that both $X$ and $D_{i}$ are smooth at a general point $x\in D_{i}$. In a neighborhood $U$ of $x$ in $X$, every $s_{j}$ is a unit if $j\neq i$. Furthermore, $D_{i}$ is defined by $t_{i}=0$ for some holomorphic function $t_{i}$ on $U$, and $s_{i}=\eta\cdot t_{i}^{r_{i}}$ for some unit holomorphic function $\eta$ on $U$. By Lemma 2.6, up to shrinking $U$, the preimage $p^{-1}(U)$ is a disjoint union of copies of $Y$, where $Y=U[\sqrt[n]{t_{i}}]$. Since $Y\to U$ is totally branched over $D_{i}\cap U$, we deduce that the ramification index of $p$ along any component of $D^{\prime}_{i}$ is equal to $n$.
@@
${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${X}$
@@
We fix a positive integer $n$ divisible by $N$. Let $p\colon V\to X$ be the natural morphism. Then the critical locus of $p$ is exactly $\Delta$. Let $H_{i}=p^{-1}(D_{i})$ for $i=1,...,m$ and let $H=\sum_{i=1}^{m}H_{i}$. Then $\mu$ is quasi-étale over $V\setminus H$. Furthermore, each $H_{i}$ is Cartier and $p^{*}D_{i}=nH_{i}$ by Lemma 2.7. The latter property implies that $\mu$ is étale over general points of every component of $H_{i}$, for $n$ is divisible by $N$. As a consequence, $\mu$ is quasi-étale.
Suppose that $\pi^{-1}(X\setminus\Delta)$ is smooth. We assume by contradiction that there is an irreducible component $Z\subseteq W_{\operatorname{\textsubscript{\rm sing}}}$ such that $\operatorname{dim}Z=\operatorname{dim}W-2$. Since $p$ is étale over $X\setminus\Delta$, we deduce that $W\to Y$ is étale over $\pi^{-1}(X\setminus\Delta)$. Therefore, $Z$ is contained in $(p\circ\mu)^{-1}(\Delta)$. Then $\mu(Z)\subseteq H$. We note that $(V,H)$ is dlt by Lemma 2.7. Since $H$ is Cartier, by (16.6.1.1) and (16.6.3.2) of [Kol92, Proposition 16.6], we deduce that $V$ is smooth around general points of $\mu(Z)$. By the Zariski’s purity theorem, we conclude that $\mu$ is étale over general points of $\mu(Z)$. Thus $W$ is smooth around general points of $Z$. This is a contradiction. ∎
@@
be a cubic polynomial in $y$ whose coefficients are holomorphic functions on $S$. Let $X\subseteq\mathbb{C}\times S$ be the closed analytic subspace defined by $F(y,s)=0$, and let $\pi\colon X\to S$ be the natural projection. Assume that, either the divisorial critical locus of $\pi$ is contained in $H$, or the critical locus is $S$. In other words, either $F(y,s)$ has three distinct roots in $y$ for any point $s\in S\setminus H$ fixed, or $F(y,s)$ has a multiple root in $y$ for a general point $s\in S$ fixed.
[…]
@@
as the pullback of $F$ on ×$\mathbb{C}\times\overline{S}$. Then there is a factorization
@@
Now we assume that the critical locus of $\pi$ is a proper subset of $S$. Let $X^{\prime}$ be the normalization of $X$ and let $\pi^{\prime}\colon X^{\prime}\to S$ be the natural morphism. Since the degree of $\pi^{\prime}$ is 3, we see that the ramification index of $\pi^{\prime}$ along any prime divisor in $X^{\prime}$ is contained in $\{1,2,3\}$. Let $\overline{X}$ be the hypersurface defined by $G$ and let $Y$ be its normalization. Then $Y$ is also the normalizaiton of $X^{\prime}\times_{S}\overline{S}$, and the natural finite morphism $Y\to\overline{S}$ is quasi-étale by Lemma 2.8. The Zariski’s purity theorem implies that it is indeed étale, for $\overline{S}$ is smooth. Since $\overline{S}$ is simply connected, we deduce that $Y$ has three connected components. Hence $\overline{X}$ has three irreducible components. The same argument as in the previous paragraph implies that each component is isomorphic to $\overline{S}$. Therefore, there is a factorization

### v6

@@ Abstract.
We prove that if $X$ is a compact complex analytic variety, which has quotient singularities in codimension 2, then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$, such that $Y$ has quotient singularities, and that the indeterminacy locus of $f^{-1}$ has codimension at least 3 in $X$.
@@ 1. Introduction
Let $X$ be a compact complex analytic variety. Assume that $X$ has quotient singularities in codimension 2. In other words, there is a closed analytic subset $V$ of codimension at least 3, such that $X\setminus V$ has quotient singularities. Then there is a projective bimeromorphic morphism $f\colon Y\rightarrow X$ such that $Y$ has quotient singularities, and that the indeterminacy locus of $f^{-1}$ has codimension at least 3 in $X$.
We note that when $X$ is projective, a stronger version of Theorem 1.2 was proved by Xu in [LT19]. The proof approximately proceeds as follows. We let $X^{\circ}$ be the largest open subset on which $X$ has quotient singularities. By considering the frame bundle on the local orbifold charts, we can obtain a smooth quasi-projective variety $Y^{\circ}$, such that $X^{\circ}=Y^{\circ}/G$, where $G=\mathrm{GL}_{n}(\mathbb{C})$ and $n$ is the dimension of $X$ (see [Kre09]). The action of $G$ on $Y^{\circ}$ only has finite stabilizers. There exists a $G$-equivariant projective compactification Y{Y} of $Y^{\circ}$, such that $G$ acts linearly on Y{Y}. Then up to applying Kirwan’s partial resolution in [Kir85], one may assume that the strictly semi-stable locus in Y{Y} is empty. Then the GIT quotient ${Y}\sslash G$ is a projective compactification of $X^{\circ}$ which has quotient singularities only. Finally, by using equivariant resolution of singularities to the birational map ${Y}\sslash G\dashrightarrow X$, we can obtain a projective compactification $Z$ of $X^{\circ}$ such that $Z$ has quotient singularities and the natural birational map $Z\dashrightarrow X$ is a morphism. The main difficulty to adapt this method in the setting of general complex analytic varieties is that, the theory of GIT relies essentially on the existence of an ample line bundle, which is not the case if $X$ is not projective.
@@ 2. Preliminaries
We fix some notation and prove some elementary results in this section. Throughout this paper, the symbol $\mathbb{D}$ stands for a disc contained in $\mathbb{C}$ centered at 0. We will write $\mathbb{D}_{r}$ if we need to stress the radius $r$ of $\mathbb{D}$. We denote the origin of $\mathbb{C}^{n}$ by $\mathbf{0}_{n}$. By a polydisc of dimension $n$, we refer to $\mathbb{D}^{n}$, which is always assumed centered at the origin of $\mathbb{C}^{n}$.
@@ 2.A. Complex analytic varieties and their singularities
A complex analytic variety $X$ is a reduced and irreducible complex analytic space. We will denote by $X_{\operatorname{\textsubscript{\rm sm}}}$ its smooth locus and by $X_{\operatorname{\textsubscript{\rm sing}}}$ its singular locus. A smooth complex analytic variety is also called a complex manifold. We say that a normal compact complex analytic variety $X$ is $\mathbb{Q}$-factorial, if for any reflexive sheaf $\mathcal{L}$ of rank one on $X$, there is some integer $m>0$ such that $(\mathcal{L}^{\otimes m})^{**}$ is locally free.
@@
[…]
$S=\mathbb{D}^{n-2}\times\{o\}\rightarrow\mathbb{D}^{n-2}\times\{\tilde{v}\}\hookrightarrow\widetilde{U}$.
Since $T$ meets $S$ transversally at $x$, we deduce that, at $\tilde{x}$, the Zariski tangent space of $\widetilde{T}$ and the Zariski tangent space of $\sigma(S)$, both viewed as the subspace of the Zariski tangent space of $\widetilde{U}$, intersect transversally. Since $\operatorname{dim}\,\widetilde{T}+\operatorname{dim}\,\sigma(S)=\operatorname{dim}\,\widetilde{U}$, and since $\widetilde{U}$ is smooth, we deduce that $\widetilde{T}$ is smooth at $\tilde{x}$. Therefore, we obtain that $\widetilde{T}$ is smooth. Furthermore, by the adjunction formula, we see that $\widetilde{T}\rightarrow T$ is the minimal resolution. This implies that $(x\in T)$ is isomorphic to $(o\in V)$ as surface singularities. ∎
@@
where $p,q$ are holomorphic functions whose terms have degrees at least 2 and 3 respectively. Replacing $z$ by zp(y,z)z-p(y,z), we can assume that
@@
Let $f\colon Y\rightarrow X$ be a proper bimeromorphic morphism between complex analytic spaces. Since we do not assume that $Y$ is normal, we say that there is a $f$-ample and $f$-exceptional divisor H-H, if $H$ is a closed analytic subspace of $Y$ contained in the $f$-exceptional locus, such that the ideal sheaf $\mathcal{I}$ of $H$ is a $f$-ample invertible sheaf (see [Nak87, Definition 1.1]). In the proof of the main theorem, we apply several local constructions around a component of the singular locus. The following lemma enables us to extend local modifications to the whole variety.
@@
Let $X$ be a complex analytic space and let $U\subseteq X$ be an open subset. Assume that there is a proper bimeromorpihc morphism $p\colon U^{\prime}\rightarrow U$ such that the indeterminacy locus $Z$ of $p^{-1}$ is compact. Then $p$ extends to some proper bimeromorpihc morphism $f\colon X^{\prime}\rightarrow X$, which is an isomorphism over $X\setminus Z$. Assume further that there is some $p$-exceptional divisor in $U^{\prime}$ which is $p$-ample over $U$, then $f$ is projective over $X$.
@@
Let $W=X\setminus Z$. Since $p$ is an isomorphism over $U\cap W$, we deduce that $U^{\prime}$ and $W$ glue together along $U\cap W$ to a complex analytic space $X^{\prime}$, and that $p$ extends to a proper bimeromorphic morphism $f\colon X^{\prime}\rightarrow X$. If there is a $p$-exceptional and $p$-ample divisor H-H, then $H$ is compact since $Z$ is compact. Thus $H$ is naturally a closed analytic subspace of $X^{\prime}$. We note that its ideal sheaf in $X$ is also invertible and $f$-ample over $X$. Hence $f\colon X^{\prime}\rightarrow X$ is projective. ∎
@@ 2.B. Finite morphisms between complex analytic varieties
In this subsection, we summarize some results concerning finite morphisms between complex analytic varieties. We recall from [GR84, page 47] that a morphism $f\colon Y\rightarrow X$ is called finite if it is closed and if every fiber is a finite set. In this case, the critical locus of $f$ is the smallest closed subset $Z\subseteq X$, such that $f$ is étale over $X\setminus Z$, that is, $f$ is locally biholomorphic on $Y\setminus f^{-1}(Z)$. By the divisorial critical locus of $f$, we refer to the codimension 1 part of $Z$. The morphism $f$ is said to be quasi-étale if its critical locus has codimension at least $2$. By abuse of notation, we say that $f$ is Galois of Galois group $G$ if its restriction over $X\setminus Z$ is Galois, and if there is an action of $G$ on $Y$, extending the Galois action of $G$ on $f^{-1}(X\setminus Z)$, such that the quotient $Y/G$ is canonically isomorphic to $X$. We refer to [Car57, Théorème 4] for the quotient of a complex analytic space by a finite group. The following theorem is due to Grauert-Remmert.
@@
Let $X$ be a complex analytic variety, and let $X^{\circ}\subseteq X$ be a dense Zariski open subset. Assume that $X^{\circ}$ is normal and we have a finite étale morphism $p\colon Y^{\circ}\rightarrow X^{\circ}$. Then $p$ extends to a finite morphism $p\colon Y\rightarrow X$ with $Y$ normal, which is unique up to isomorphism.
@@
Let $r\colon X^{\prime}\rightarrow X$ be the normalization. By assumption, $r$ is an isomorphism over $X^{\circ}$. Hence up to replacing $X$ by $X^{\prime}$, we may assume that $X$ is normal. In this case, the theorem is proved in [GR71, Théorème XII.5.4]. ∎
@@
$X[\sqrt[k_{1}]{s_{1}},...,\sqrt[k_{m}]{s_{m}}]=\mathrm{Spec}_{\mathcal{O_{X}}}\,\mathcal{A}$
@@
With the notation above, let $x\in X$ be a point and let $s^{\prime}_{1},...,s^{\prime}_{m}$ be holomorphic functions on $X$. Assume that there are unit functions $\mu_{1},...,\mu_{m}$ such that $s^{\prime}_{i}=\mu_{i}\cdot s_{i}$ for all i=1,,mi=1,...,m. Then, up to shrinking $X$ around $x$, there is an isomorphism over $X$
$X[\sqrt[k_{1}]{s_{1}},...,\sqrt[k_{m}]{s_{m}}]\cong X[\sqrt[k_{1}]{s^{\prime}_{1}},...,\sqrt[k_{m}]{s^{\prime}_{m}}].$
In particular, if $x^{\prime}\in X[\sqrt[k_{1}]{s_{1}},...,\sqrt[k_{m}]{s_{m}}]$ is a point lying over $x$, then the germ $(x^{\prime}\in X[\sqrt[k_{1}]{s_{1}},...,\sqrt[k_{m}]{s_{m}}])$ depends only on the integers $k_{1},...,k_{m}$, and the stalks $(\mathcal{I}_{1})_{x},...,(\mathcal{I}_{m})_{x}$ at $x$ of the invertible ideal sheaves generated by $s_{1},...,s_{m}$ respectively.
@@
With the notation above, we assume that $X$ is a normal complex analytic variety. Suppose that, for i=1,,mi=1,...,m, there is an integer $r_{i}>0$ and a divisor $D_{i}$ such that $r_{i}D_{i}$ is defined by $s_{i}=0$. Then for any integer $n>0$, we can construct the cyclic cover $X[\sqrt[nr_{1}]{s_{1}},...,\sqrt[nr_{m}]{s_{m}}]$ over $X$. Let
$X[\sqrt[nr_{1}]{s_{1}},...,\sqrt[nr_{m}]{s_{m}}]^{nor}$
be its normalization and let $p\colon X[\sqrt[nr_{1}]{s_{1}},...,\sqrt[nr_{m}]{s_{m}}]^{nor}\rightarrow X$ be the natural morphism.
@@
With the notation above, we let $x\in X$ be a point and let $U\subseteq X$ be an open neighborhood of $x$. Assume that for each i=1,,mi=1,...,m, the divisor $l_{i}D_{i}|_{U}$ is defined by $t_{i}=0$, where $l_{i}>0$ is an integer and $t_{i}$ is a holomorphic function on $U$. Then, up to shrinking $U$, every connected component of $U[\sqrt[nl_{1}]{t_{1}},...,\sqrt[nl_{m}]{t_{m}}]^{nor}$ is isomorphic to every connected component of $p^{-1}(U)$.
@@
With the notation above, we set $X^{\prime}=X[\sqrt[nr_{1}]{s_{1}},...,\sqrt[nr_{m}]{s_{m}}]^{nor}$. Assume that every $D_{i}$ is reduced and irreducible. Then the following properties hold.
@@
For the item (1), we first prove the statement on the ramification indices of $p$. Since $X$ is normal and $D_{i}$ is reduced, we deduce that both $X$ and $D_{i}$ are smooth at a general point $x\in D_{i}$. In a neighborhood $U$ of $x$ in $X$, every $s_{j}$ is a unit if $j\neq i$. Furthermore, $D_{i}$ is defined by $t_{i}=0$ for some holomorphic function $t_{i}$ on $U$, and $s_{i}=\eta\cdot t_{i}^{r_{i}}$ for some unit holomorphic function $\eta$ on $U$. By Lemma 2.6, up to shrinking $U$, the preimage $p^{-1}(U)$ is a disjoint union of copies of $Y$, where $Y=U[\sqrt[n]{t_{i}}]$. Since $Y\rightarrow U$ is totally branched over $D_{i}\cap U$, we deduce that the ramification index of $p$ along any component of $D^{\prime}_{i}$ is equal to $n$.
@@
Let $(X,\Delta)$ be a reduced dlt pair. Let $D_{1},...,D_{m}$ be the irreducible components of $\Delta$. Assume that $r_{i}D_{i}$ is defined by $s_{i}=0$ for some holomorphic function $s_{i}$ on $X$ and for some integer $r_{i}>0$. Let $\pi\colon Y\rightarrow X$ be a finite morphism with $Y$ normal, whose divisorial critical locus is contained in $\Delta$. Let $N$ be the lcm of all the coefficients in $\pi^{*}D_{i}$ for i=1,,mi=1,...,m. Then for any positive integer $n$ divisible by $N$, the following properties hold. Let $V=X[\sqrt[nr_{1}]{s_{1}},...,\sqrt[nr_{m}]{s_{m}}]^{nor}$ and let $W$ be the normalization of $Y\times_{X}V$. Then the natural morphism $\mu\colon W\rightarrow V$ is quasi-étale. Moreover, if $\pi^{-1}(X\setminus\Delta)$ is smooth, then $W$ is smooth in codimension 2.
@@
We fix a positive integer $n$ divisible by $N$. Let $p\colon V\rightarrow X$ be the natural morphism. Then the critical locus of $p$ is exactly $\Delta$. Let $H_{i}=p^{-1}(D_{i})$ for i=1,,mi=1,...,m and let $H=\sum_{i=1}^{m}H_{i}$. Then $\mu$ is quasi-étale over $V\setminus H$. Furthermore, each $H_{i}$ is Cartier and $p^{*}D_{i}=nH_{i}$ by Lemma 2.7. The latter property implies that $\mu$ is étale over general points of every component of $H_{i}$, for $n$ is divisible by $N$. As a consequence, $\mu$ is quasi-étale.
Suppose that $\pi^{-1}(X\setminus\Delta)$ is smooth. We assume by contradiction that there is an irreducible component $Z\subseteq W_{\operatorname{\textsubscript{\rm sing}}}$ such that $\operatorname{dim}Z=\operatorname{dim}W-2$. Since $p$ is étale over $X\setminus\Delta$, we deduce that $W\rightarrow Y$ is étale over $\pi^{-1}(X\setminus\Delta)$. Therefore, $Z$ is contained in $(p\circ\mu)^{-1}(\Delta)$. Then $\mu(Z)\subseteq H$. We note that $(V,H)$ is dlt by Lemma 2.7. Since $H$ is Cartier, by (16.6.1.1) and (16.6.3.2) of [Kol92, Proposition 16.6], we deduce that $V$ is smooth around general points of $\mu(Z)$. By the Zariski’s purity theorem, we conclude that $\mu$ is étale over general points of $\mu(Z)$. Thus $W$ is smooth around general points of $Z$. This is a contradiction. ∎
@@
be a cubic polynomial in $y$ whose coefficients are holomorphic functions on $S$. Let $X\subseteq\mathbb{C}\times S$ be the closed analytic subspace defined by $F(y,s)=0$, and let $\pi\colon X\rightarrow S$ be the natural projection. Assume that, either the divisorial critical locus of $\pi$ is contained in $H$, or the critical locus is $S$. In other words, either $F(y,s)$ has three distinct roots in $y$ for any point $s\in S\setminus H$ fixed, or $F(y,s)$ has a multiple root in $y$ for a general point $s\in S$ fixed.
Let $\overline{S}=S[\sqrt[6]{T_{1}},...,\sqrt[6]{T_{n}}]$ and let $\mu\colon\overline{S}\rightarrow S$ be the natural morphism. We define
@@
as the pullback of $F$ on $\mathbb{C}\times\overline{S}$. Then there is a factorization
@@
Now we assume that the critical locus of $\pi$ is a proper subset of $S$. Let $X^{\prime}$ be the normalization of $X$ and let $\pi^{\prime}\colon X^{\prime}\rightarrow S$ be the natural morphism. Since the degree of $\pi^{\prime}$ is 3, we see that the ramification index of $\pi^{\prime}$ along any prime divisor in $X^{\prime}$ is contained in $\{1,2,3\}$. Let $\overline{X}$ be the hypersurface defined by $G$ and let $Y$ be its normalization. Then $Y$ is also the normalizaiton of $X^{\prime}\times_{S}\overline{S}$, and the natural finite morphism $Y\rightarrow\overline{S}$ is quasi-étale by Lemma 2.8. The Zariski’s purity theorem implies that it is indeed étale, for $\overline{S}$ is smooth. Since $\overline{S}$ is simply connected, we deduce that $Y$ has three connected components. Hence $\overline{X}$ has three irreducible components. The same argument as in the previous paragraph implies that each component is isomorphic to $\overline{S}$. Therefore, there is a factorization